OpenStudy (anonymous):

Let f(x) = 5x + 12. Find f^-1(x). I got x-2/5 x minus 2 over the quantity 5.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hello. o:

OpenStudy (misty1212):

function says ''multiply by 5, add 12"

OpenStudy (misty1212):

inverse says "subtract 12, divide by 5"

OpenStudy (anonymous):

Supposed to do the inverse, so I did. Lemme screenshot.

OpenStudy (misty1212):

you wrote \[\frac{x-2}{5}\] but i bet you meant \[\frac{x-12}{5}\] since there is not 2 in it

OpenStudy (misty1212):

no i got it you got it too i think

OpenStudy (anonymous):

Yeah I got that. xD I forgot the 1 to make it 12. Here's le work. http://prntscr.com/8ysskd

OpenStudy (anonymous):

I wrote two and from there I thought it was, derp. xD

OpenStudy (misty1212):

man is that a bad mistake (not yours, the screen shot)

OpenStudy (misty1212):

where did that come from?

OpenStudy (anonymous):

I used a program called lightshot to take screenshots. Helpful little thing, really.

OpenStudy (misty1212):

lol no dear i dont mean the screenshot itself i mean who wrote it?

OpenStudy (misty1212):

does it come from a text? your teacher? on line system?

OpenStudy (anonymous):

Oh, someone on Brainly. I wrote the same thing, oh wow that's the wrong screenshot. XD That was for another question I worked out. I like making sure I do everything perfectly because math is my weakpoint. When I come here it's to make sure I understand everything to its fullest. Right now functions are my weakest point, but I work them out on my own then come here to see if I'm right.

OpenStudy (anonymous):

Most people come here for answers, I rather be taught so I can do it on my own. I can't trust any answers I find on the web, they may be wrong, so if I learn it by someone here, then woohoo, I can make sure I got it down pat.

OpenStudy (misty1212):

ok i don't mean to be critical but there is a huge mistake there

OpenStudy (anonymous):

Where? The screenshot?

OpenStudy (misty1212):

\[f(x)=x^2-81\] is not a one to one function it is a parabola and doesn't pass the horizontal line test for example \[f(10)=-19=f(-10)\]

OpenStudy (anonymous):

I got \[\frac{ x-12 }{ 5 }\]

OpenStudy (misty1212):

when you solve you should get \[\pm\sqrt{x+81}\]

OpenStudy (misty1212):

yeah that is right, i am talking about the screen shot

OpenStudy (anonymous):

That's what I got. It should have two answers because of the square root thing so the = and - are there. Where's the mistake?

OpenStudy (anonymous):

+ and -*

OpenStudy (misty1212):

i didn't see it in the screen shot is all

OpenStudy (anonymous):

Ah. Thanks!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]